### v1

@@ 6. Uniqueness of the single-fold principal train track map
for $k\in\{2,\dots n-1\}$.
The final two recursively defined edges $\alpha_{n-2}=[v_{2},v_{3}]$ and $\alpha_{n-1}=[v_{1},v_{2}]$ both contain $v_{2}$. Moreover, $e_{1}=[v_{0},v_{2}]$ and $e_{3}=[v_{2},v_{4}]$ contain $v_{2}$. Since $v_{0},v_{1},v_{3},$ and $v_{4}$ are distinct vertices, all four of these edges are distinct. Therefore $v_{2}$ is contained in at least 4 distinct edges in $\Gamma$. However, this contradicts that $v_{2}$ should have valence 3. Thus for $r\geq 4$ can there exist a single-fold irreducible tt map $h$ representing a principal fully irreducible element of $\textup{Out}(F_{r})$.
Figure 3. This figure depicts the Rank-3 Principal Stratum Automaton. Permutations of the automata are included in green. Compositions of included permutations are implicitly included.

### v2

@@ 6. Uniqueness of the single-fold principal train track map
for each $k\in\{2,\dots n-1\}$.
The final two recursively defined edges $\alpha_{n-2}=[v_{2},v_{3}]$ and $\alpha_{n-1}=[v_{1},v_{2}]$ both contain $v_{2}$. Moreover, $e_{1}=[v_{0},v_{2}]$ and $e_{3}=[v_{2},v_{4}]$ contain $v_{2}$. Since $v_{0},v_{1},v_{3},$ and $v_{4}$ are distinct vertices, all four of these edges are distinct. Therefore $v_{2}$ is contained in at least 4 distinct edges in $\Gamma$. This contradicts that $v_{2}$ should have valence 3. Thus for each $r\geq 4$ there cannot exist a single-fold irreducible tt map $h$ representing a principal fully irreducible element of $\textup{Out}(F_{r})$. ∎
Figure 3. This figure depicts the Rank-3 Principal Stratum Automaton. Permutations of the automata are in green; note that if moving right to left across the page one must read the permuatation right to left as well. Compositions of included permutations, as well as color-preserving graph symmetries, are implicitly included. (Symmetries are included in blue at II as an example.) The map of Figure 2 is in dotted gold.